# Visual Sparse Steering: Improving Zero-shot Image Classification with Sparsity Guided Steering Vectors

### v1

@@ 3.2.1 VS2: Visual Sparse Steering
Learning latent concept space. Given an input image $X\in\mathcal{X}$ and its corresponding embedding ${\bm{x}}\in{\mathbb{R}}^{d}$ extracted from layer $i$ of a Vision Foundation Model (VFM), we define a Concept Encoder $E_{c}\colon{\mathbb{R}}^{d}\to{\mathbb{R}}^{|{\mathbb{C}}|}$ that maps the embedding ${\bm{x}}$ into a sparse concept activation vector:
${\bm{c}}=E_{c}({\bm{x}})=(c_{1},\dots,c_{|{\mathbb{C}}|}).$ (5)
By imposing sparsity constraints described in Sec. 3.1 on the concept activations, the encoder $E_{c}$ is encouraged to identify disentangled, semantically meaningful concepts. To map the concepts into the original representation space, we define a Concept Decoder $D_{c}\colon{\mathbb{R}}^{|{\mathbb{C}}|}\to{\mathbb{R}}^{d}$ that reconstructs the original embedding from these concepts:
$\tilde{{\bm{x}}}=D_{c}({\bm{c}}).$ (6)
Since explicit supervision for concept identification is typically unavailable, we train the encoder-decoder pair jointly on training data $\mathcal{X}$, enforcing both sparsity and accurate embedding reconstruction to implicitly uncover meaningful concepts.
@@
Constructing steering vector. At inference stage, emerged sparse concepts are deemed semantically significant and critical for downstream tasks, and we enhance them via a Concept Upweighting Function $U:{\mathbb{R}}^{|{\mathbb{C}}|}\to{\mathbb{R}}^{|{\mathbb{C}}|}$. In implementation, we simply scale the original concept activations ${\bm{c}}$ by a factor of $\gamma$ to amplify their effects:
${\bm{c}}^{\prime}=U({\bm{c}})=\gamma\times{\bm{c}}.$ (7)
Reconstructing this modified concept vector ${\bm{c}}^{\prime}$ provides a conceptually steered embedding:
$\tilde{{\bm{x}}}^{\prime}=D_{c}({\bm{c}}^{\prime}).$ (8)
Since reconstruction steps inherently introduce approximation errors, we define the steering vector ${\bm{v}}$ as:
${\bm{v}}=\tilde{{\bm{x}}}^{\prime}-\tilde{{\bm{x}}},$ (9)
which explicitly represents the semantic shift induced by upweighting and mitigates the effects of the reconstruction errors.
Steer the target emebedding. Obtaining the steering vector ${\bm{v}}$ that aims to amplify the latent concepts embedded in ${\bm{x}}$, we shift the original embedding by adding the steering vector back as:
$\hat{{\bm{x}}}={\bm{x}}+\lambda\,{\bm{v}}.$ (10)
where $\lambda$ is a hyperparameter that controls the magnitude of steering. To stabilize the representation, we further rescale the $\ell_{2}$ norm of the steered embedding to its original scale:
$\hat{{\bm{x}}}=\frac{\hat{{\bm{x}}}\cdot\|{\bm{x}}\|_{2}}{\|\hat{{\bm{x}}}\|_{2}}.$ (11)
Intuitively, this procedure precisely moves the embedding ${\bm{x}}$ along directions defined by semantically relevant, interpretable concepts, thus improving interpretability and enhancing robustness against spurious correlations. We refer Fig. 1 (upper) for a schematic overview of VS2.
3.2.2 VS2++: Visual Sparse Steering with Cached Data
In VS2, we leverage the learned SAEs to implicitly decide the critical concepts for a given latent embedding, and treat all identified concepts equally important. One natural speculation is that there are certain concepts that are unique and more important to the target underlying distribution, while others are more general even spuriously correlated. To this end, we propose VS2++ that allows the access of cached training embedding at test-time to enable a more nuanced control over the concept steering.
@@ 3.2.2 VS2++: Visual Sparse Steering with Cached Data
Constructing Contrastive Groups. Given a query embedding ${\bm{x}}_{q}$ and a cached dataset of embeddings $\{{\bm{x}}_{i}\}_{i=1}^{M}$, VS2++ first retrieve the set of $N$ nearest neighbors $\mathcal{N}_{N}({\bm{x}}_{q})$ for a query embedding ${\bm{x}}_{q}$ is defined as:
@@
where $S(,)$ denotes the measurement of cosine similarity.
After retrieving $\mathcal{N}_{N}({\bm{x}}_{q})$, we obtain the pseudo-label for the query embedding $\hat{y}_{q}\;=\;\arg\max_{y}\,P(y\,\mid\,{\bm{x}}_{q})$, where $P(y\,\mid\,{\bm{x}})$ is the classifier’s predicted probability for label $y$. Similarly, for each neighbor ${\bm{x}}_{i}\in\mathcal{N}_{N}({\bm{x}}_{q})$, we compute its pseudo-label: $\hat{y}_{i}\;=\;\arg\max_{y}\,P(y\,\mid\,{\bm{x}}_{i})$.
We then construct the positive group, ${\mathbb{S}}^{+}({\bm{x}}_{q})$, by selecting neighbors sharing the same pseudo-label as the query embedding:
${\mathbb{S}}^{+}({\bm{x}}_{q})\;=\;\{{\bm{x}}_{i}\,\in\,\mathcal{N}_{N}({\bm{x}}_{q})\;\mid\;\hat{y}_{i}=\hat{y}_{q}\}.$ (13)
Likewise, the negative group, ${\mathbb{S}}^{-}({\bm{x}}_{q})$, is formed by the remaining neighbors:
${\mathbb{S}}^{-}({\bm{x}}_{q})\;=\;\mathcal{N}_{N}({\bm{x}}_{q})\;\setminus\;{\mathbb{S}}^{+}({\bm{x}}_{q}).$ (14)
Selective Steering. For each embedding ${\bm{x}}_{i}\in{\mathbb{S}}^{+}({\bm{x}}_{q})$, VS2++ first applies VS2 (Eq. 7,8,9) to acquire a directional vector ${\bm{v}}_{i}^{p}$ that underscores the underlying concepts within the given embedding. We repeat the same procedure for each ${\bm{x}}_{j}\in{\mathbb{S}}^{-}({\bm{x}}_{q})$ to obtain the negative steering vectors ${\bm{v}}_{j}^{n}$. We then average these per-embedding steering vectors within each group, yielding the positive anchor
@@
Finally, we obtain the contrastive steering vector by subtracting the negative anchor from the positive anchor:
${\bm{v}}\;=\;\bar{{\bm{v}}}^{p}\;-\;\bar{{\bm{v}}}^{n},$ (17)
and apply it to the query embedding ${\bm{x}}_{q}$ in the same manner as VS2:
@@
This selective scheme thus anchors the steering directions in concepts consistently activated within the positive group, while suppressing undesired features highlighted by the negative group. Since VS2++ retrieves N-nearest embeddings of a query to construct both positive and negative groups, negative group contains non-trivial hard cases.
@@ 4 Experiments
CIFAR-100 CUB-200 Tiny-IN Method ViT-B/32 ViT-B/16 ViT-B/32 ViT-B/16 ViT-B/32 ViT-B/16 Zero-shot (no retrieval) $\text{CLIP}_{\text{ZS}}$ 61.07 ( 0 ) 63.96 ( 0 ) 51.76 ( 0 ) 55.06 ( 0 ) 56.64 ( 0 ) 61.08 ( 0 ) $\text{SAE}_{\text{REC}}^{\text{A}}$ 58.01 ( –3.06 ) 64.05 ( +0.09 ) 47.45 ( -4.31 ) 51.81 ( -3.25 ) 30.56 ( –26.08 ) 52.96 ( -8.12 ) $\text{SAE}_{\text{REC}}^{\text{F}}$ 58.22 ( –2.85 ) 63.42 ( –0.54 ) 48.08 ( -3.68 ) 51.43 ( -3.63 ) 36.33 ( –20.31 ) 54.84 ( -6.24 ) $\text{SAE}_{\text{REC}}^{\text{F}+\gamma}$ 62.69 ( +1.62 ) 66.81 ( +2.85 ) 49.41 ( -2.35 ) 53.28 ( -1.78 ) 39.49 ( –17.15 ) 58.81 ( -2.27 ) VS2 (ours) 64.52 ( +3.45 ) 68.08 ( +4.12 ) 52.69 ( +0.93 ) 56.14 ( +1.08 ) 58.14 ( +1.50 ) 62.92 ( +1.84 ) VS2 + PASS (ours) 70.64 ( +9.57 ) 68.23 ( +4.27 ) 52.97 ( +1.21 ) 56.63 ( +1.57 ) 57.94 ( +1.30 ) 62.98 ( +1.90 ) RAG-enhanced (oracle cache) Weighted RAG 76.43 ( +15.36 ) 69.78 ( +5.82 ) 58.32 ( +6.56 ) 60.65 ( +5.59 ) 72.53 ( +15.89 ) 67.84 ( +6.75 ) CLIP Steering Vector 81.85 ( +20.78 ) 84.12 ( +20.16 ) 56.42 ( +4.66 ) 61.51 ( +6.45 ) 80.38 ( +23.74 ) 84.07 ( +22.99 ) VS2++ 81.95 ( +20.88 ) 85.40 ( +21.44 ) 58.84 ( +7.08 ) 61.91 ( +6.85 ) 73.27 ( +16.63 ) 81.55 ( +20.47 ) RAG-enhanced (non-oracle cache) CLIP Steering Vector 77.22 ( +16.15 ) 78.97 ( +15.01 ) 47.89 ( -3.87 ) 53.95 ( -1.11 ) 74.06 ( +17.42 ) 76.79 ( +15.71 ) VS2++ 77.11 ( +16.04 ) 79.14 ( +15.18 ) 52.81 ( +1.05 ) 57.02 ( +1.96 ) 72.12 ( +15.48 ) 76.92 ( +15.84 )
Experimental Setup.
We test VS2 and VS2++ on CIFAR-100 krizhevsky2009learning , Tiny-ImageNet le2015tiny , and CUB-200 wah2011caltech three datasets, covering standard, complex, and fine-grained classification tasks. These datasets vary in scale, diversity, and class distribution, allowing us to assess VS2 and VS2++ in both general and imbalanced classification scenarios. For vision foundation model, we use CLIP radford2021learning with ViT-B/32, and ViT-B/16 backbones. We train top-$k$ Sparse Autoencoders (SAEs) on the CLS token embeddings extracted from every layer of CLIP’s vision transformer, using the training data of each dataset to learn sparse latent representations. More details regarding the training of the top-$k$ Sparse Autoencoders can be found in the Appendix C.
4.1 Visual Sparse Steering methods
4.1.1 VS2: Visual Sparse Steering without Contrastive Data
@@
We evaluate the effectiveness of VS2 in the absence of external contrastive data. As baselines, we report: (1) the zero-shot performance of the original CLIP model ($\text{CLIP}_{\text{ZS}}$); (2) $\text{CLIP}_{\text{REC}}^{\text{F}}$, which uses the SAE-reconstructed CLS token from the final layer without any steering; and (3) $\text{CLIP}_{\text{REC}}^{\text{A}}$, which uses the reconstructed CLS tokens from all layers. These baselines isolate the effect of reconstruction alone, without applying any modification or steering vector. All top-$k$ SAEs are trained on CLS token representations across all transformer layers. In addition to these baselines, we report $\text{SAE}_{\text{REC}}^{\text{F}+\gamma}$ that applies steering by upweighting the top-$k$ sparse features of the reconstructed CLS token of the final layer with a scaling factor of $\gamma=1.5$. This variant isolates the impact of steering the SAE’s latent activations by amplifying their top-$k$ before the CLS embedding is reconstructed, letting us quantify the benefit attributable to this targeted up-weighting alone. In Appendix G, we provide the pseudocode for main differences between the proposed baselines and VS2. In Appendix H, we also provide comparative results of our method compared to other baselines such as Splice (bhalla2024interpreting, ). Note that Splice uses an external vocabulary to construct latent features; in contrast, we use unsupervised SAEs to learn latent features for VS2 and VS2++.
@@ Results.
From Table 1, we demonstrate that using only SAE reconstructions generally reduces zero-shot performance in all datasets; we hypothesize this is due to reconstruction errors in SAEs (engels2025decomposing, ). However, when SAE features are amplified in the reconstruction, performance can improve, as demonstrated by $\text{SAE}_{\text{REC}}^{\text{F}+\gamma}$ for CIFAR-100. Specifically, $\text{SAE}_{\text{REC}}^{\text{F}+\gamma}$ outperforms the zero-shot baseline on CIFAR-100 by 1.62% and 2.85% using the ViT-B/32 and ViT-B/16 backbones respectively. However, $\text{SAE}_{\text{REC}}^{\text{F}+\gamma}$ is worse than the CLIP baseline on CUB-200 and Tiny-ImageNet. In contrast, our method VS2 consistently outperforms all baseline methods across all datasets and ViT backbones, surpassing zero-shot CLIP by 3.45% and 4.12% on CIFAR-100, 0.93% and 1.08% on CUB-200, and 1.50% and 1.84% on Tiny-ImageNet with ViT-B/32 and ViT-B/16 backbones, respectively. In summary, we demonstrate that our proposed steering vectors using SAEs can improve downstream classification performance on vision tasks without relying on any external data or supervision. In Appendix A, we also provide evidence that the sparse features are interpretable; they can capture subtle class-specific attributes and concepts such as birds with gray upperpart, white underpart, white-colored throat and black eyes. Finally, in Appendix B, we provide a sensitivity analysis of VS2 hyperparameters $\gamma$ and $\lambda$ showing the robustness of our proposed methods.
4.1.2 VS2++: Visual Sparse Steering with Selectively Amplifying Features
In VS2, we leverage learned SAEs to identify critical concepts from a latent embedding, treating all activated concepts as equally important. However, some concepts may be more discriminative or unique to the target class, while others may be generic or even spuriously correlated with a class label. To enable finer control over concept importance, we introduce VS2++, which accesses cached training data at test time to guide steering more selectively. As a proxy for identifying salient features, we adopt a Retrieval-Augmented Generation (RAG) approach: we retrieve the top-$50$ most similar training images to the test input using DINOv2 oquab2023dinov2 and average their sparse activations. To construct the positive and negative sets as shown in equations 13, 14, we assign pseudolabels using a CLIP classifier; images with the most frequent pseudolabel form the positive set, while the remaining images serve as negatives. This setup corresponds to the Non-oracle cache in Table 1. For comparison, we also report an unrealistic scenario that serves as an upper-bound i.e. Oracle cache setting that uses ground-truth labels to define positives and negatives.
Baselines.
As a baseline in the Retrieval-Augmented Generation (RAG) pipeline lewis2020retrieval , we combine the test query embedding with a weighted aggregation of retrieved document embeddings. Additional details on the weighting strategy are provided in Appendix D. We also include a non-SAE variant that constructs a steering vector by averaging the CLIP embeddings of the positive and negative sets and taking their difference, yielding a purely contrastive direction without relying on Sparse Autoencoders.

### v2

@@ 3.1.1 VS2: Visual Sparse Steering
Learning latent concept space. Given an input image $X\in\mathcal{X}$ and its corresponding embedding ${\bm{x}}\in{\mathbb{R}}^{d}$ extracted from layer $i$ of a Vision Foundation Model (VFM), we define a Concept Encoder $E_{c}\colon{\mathbb{R}}^{d}\to{\mathbb{R}}^{|{\mathbb{C}}|}$ that maps the embedding $\bm{x}$} into a sparse feature activation vector: ${\bm{c}}=E_{c}({\bm{x}})=(c_{1},\dots,c_{|{\mathbb{C}}|})$. By imposing sparsity constraints on the feature activations, the encoder $E_{c}$ is encouraged to identify disentangled, semantically meaningful features. To map the features into the original representation space, we define a Concept Decoder $D_{c}\colon{\mathbb{R}}^{|{\mathbb{C}}|}\to{\mathbb{R}}^{d}$ that reconstructs the original embedding from these features: $\tilde{{\bm{x}}}=D_{c}({\bm{c}})$. Since explicit supervision for feature identification is typically unavailable, we train the encoder-decoder pair jointly on training data $\mathcal{X}$, enforcing both sparsity and accurate embedding reconstruction to implicitly uncover meaningful features.
@@
Constructing steering vector. At inference time, emerged sparse features are assumed semantically significant and critical for downstream tasks, and we enhance them via a Concept Upweighting Function $U:{\mathbb{R}}^{|{\mathbb{C}}|}\to{\mathbb{R}}^{|{\mathbb{C}}|}$. In implementation, we simply scale the original feature activations $\bm{c}$} by a factor of $\gamma$ to amplify their effects: ${\bm{c}}^{\prime}=U({\bm{c}})=\gamma\times{\bm{c}}$. Reconstructing this modified feature vector $\bm{c}$$^{\prime}$ provides a conceptually steered embedding: $\tilde{{\bm{x}}}^{\prime}=D_{c}({\bm{c}}^{\prime})$. Since reconstruction steps inherently introduce approximation errors, we define the steering vector $\bm{v}$} as: ${\bm{v}}=\tilde{{\bm{x}}}^{\prime}-\tilde{{\bm{x}}}$. which explicitly represents the semantic shift induced by upweighting and aims to mitigate the effects of the reconstruction errors.
Steer the target embedding. Obtaining the steering vector $\bm{v}$} that aims to amplify the latent features embedded in $\bm{x}$}, we shift the original embedding by adding the steering vector back as: $\hat{{\bm{x}}}={\bm{x}}+\lambda\,{\bm{v}}$. where $\lambda$ is a hyperparameter that controls the magnitude of steering. To stabilize the representation, we further rescale the $\ell_{2}$ norm of the steered embedding to its original scale: $\hat{{\bm{x}}}=\frac{\hat{{\bm{x}}}\cdot\|{\bm{x}}\|_{2}}{\|\hat{{\bm{x}}}\|_{2}}.$ Intuitively, this procedure precisely moves the embedding $\bm{x}$} along directions defined by semantically relevant, salient features, thus enhancing robustness against spurious correlations. We refer Fig. 1 (left) for a schematic overview of VS2.
3.1.2 VS2++: Visual Sparse Steering with Unlabeled Data
In VS2, we leverage the learned SAEs to identify the most salient features for a given latent representation, but treat all identified features equally important by equally amplifying them. In vision models, there are potentially many redundant features, e.g., when predicting labels, we want to ignore background features such as the sky. To selectively amplify only important features more robustly and non-uniformly, we relax the problem, allowing the existence of an unlabeled external vision corpus, and propose VS2++ that leverages this additional unlabeled data to construct steering vectors. For a given target image, VS2++ retrieves the Top-N most similar visual embeddings and constructs positive and negative groups according to generated pseudo labels. The VS2++ Steering Vector is then constructed via subtracting the average SAE features of the negative group from the positive group. Intuitively, VS2++ amplifies the SAE features most relevant for distinguishing the positive and negative groups.
@@ 3.1.2 VS2++: Visual Sparse Steering with Unlabeled Data
Constructing Contrastive Groups. Given a query embedding ${\bm{x}}_{q}$ and an unlabeled dataset of embeddings $\{{\bm{x}}_{i}\}_{i=1}^{M}$, VS2++ first retrieve the set of $N$ nearest neighbors $\mathcal{N}_{N}({\bm{x}}_{q})$ for a query embedding ${\bm{x}}_{q}$ is defined as:
@@
where $S(,)$ denotes the measurement of cosine similarity. After retrieving $\mathcal{N}_{N}({\bm{x}}_{q})$, we obtain the pseudo-label for the query embedding $\hat{y}_{q}\;=\;\arg\max_{y}\,P(y\,\mid\,{\bm{x}}_{q})$, where $P(y\,\mid\,{\bm{x}})$ is the classifier’s predicted probability for label $y$. Similarly, for each neighbor ${\bm{x}}_{i}\in\mathcal{N}_{N}({\bm{x}}_{q})$, we compute its pseudo-label: $\hat{y}_{i}\;=\;\arg\max_{y}\,P(y\,\mid\,{\bm{x}}_{i})$. We then construct the positive group, ${\mathbb{S}}^{+}({\bm{x}}_{q})$, by selecting neighbors sharing the same pseudo-label as the query embedding: ${\mathbb{S}}^{+}({\bm{x}}_{q})\;=\;\{{\bm{x}}_{i}\,\in\,\mathcal{N}_{N}({\bm{x}}_{q})\;\mid\;\hat{y}_{i}=\hat{y}_{q}\}.$ Likewise, the negative group, ${\mathbb{S}}^{-}({\bm{x}}_{q})$, is formed by the remaining neighbors: ${\mathbb{S}}^{-}({\bm{x}}_{q})\;=\;\mathcal{N}_{N}({\bm{x}}_{q})\;\setminus\;{\mathbb{S}}^{+}({\bm{x}}_{q}).$
Selective Steering. For each embedding ${\bm{x}}_{i}\in{\mathbb{S}}^{+}({\bm{x}}_{q})$, VS2++ first obtains the steering vectors as in VS2 acquiring a directional vector ${\bm{v}}_{i}^{p}$ that underscores the underlying features within the given embedding. We repeat the same procedure for each ${\bm{x}}_{j}\in{\mathbb{S}}^{-}({\bm{x}}_{q})$ to obtain the negative steering vectors ${\bm{v}}_{j}^{n}$. We then average these per-embedding steering vectors within each group, yielding the positive anchor
@@
Finally, we obtain the contrastive steering vector by subtracting the negative anchor from the positive anchor: ${\bm{v}}\;=\;\bar{{\bm{v}}}^{p}\;-\;\bar{{\bm{v}}}^{n},$ and apply it to the query embedding ${\bm{x}}_{q}$ in the same manner as VS2:
@@
This selective scheme thus anchors the steering directions in features consistently activated within the positive group, while suppressing undesired features highlighted by the negative group. Since VS2++ retrieves N-nearest embeddings of a query to construct both positive and negative groups, negative group contains non-trivial hard cases. We refer Fig. 1 (right) for a schematic overview of VS2++.
@@ 4 Experiments
CIFAR-100 CUB-200 Tiny-IN Method ViT-B/32 ViT-B/16 ViT-B/32 ViT-B/16 ViT-B/32 ViT-B/16 Label-free, unsupervised steering $\text{CLIP}_{\text{ZS}}$ 61.07 ( 0 ) 63.96 ( 0 ) 51.76 ( 0 ) 55.06 ( 0 ) 56.64 ( 0 ) 61.08 ( 0 ) $\text{SAE}_{\text{REC}}^{\text{A}}$ 58.01 ( -3.06 ) 64.05 ( +0.09 ) 47.45 ( -4.31 ) 51.81 ( -3.25 ) 30.56 ( -26.08 ) 52.96 ( -8.12 ) $\text{SAE}_{\text{REC}}^{\text{F}}$ 58.22 ( -2.85 ) 63.42 ( -0.54 ) 48.08 ( -3.68 ) 51.43 ( -3.63 ) 36.33 ( -20.31 ) 54.84 ( -6.24 ) $\text{SAE}_{\text{REC}}^{\text{F}+\gamma}$ 62.69 ( +1.62 ) 66.81 ( +2.85 ) 49.41 ( -2.35 ) 53.28 ( -1.78 ) 39.49 ( -17.15 ) 58.81 ( -2.27 ) VS2 (ours) 64.52 ( +3.45 ) 68.08 ( +4.12 ) 52.69 ( +0.93 ) 56.14 ( +1.08 ) 58.14 ( +1.50 ) 62.92 ( +1.84 ) RAG-enhanced (oracle unlabeled external data) Weighted RAG 76.43 ( +15.36 ) 69.78 ( +5.82 ) 58.32 ( +6.56 ) 60.65 ( +5.59 ) 72.53 ( +15.89 ) 67.84 ( +6.75 ) CLIP Steering Vector 81.85 ( +20.78 ) 84.12 ( +20.16 ) 56.42 ( +4.66 ) 61.51 ( +6.45 ) 80.38 ( +23.74 ) 84.07 ( +22.99 ) VS2++ 81.95 ( +20.88 ) 85.40 ( +21.44 ) 58.84 ( +7.08 ) 61.91 ( +6.85 ) 73.27 ( +16.63 ) 81.55 ( +20.47 ) RAG-enhanced (non-oracle unlabeled external data) CLIP Steering Vector 77.22 ( +16.15 ) 78.97 ( +15.01 ) 47.89 ( -3.87 ) 53.95 ( -1.11 ) 74.06 ( +17.42 ) 76.79 ( +15.71 ) VS2++ 77.11 ( +16.04 ) 79.14 ( +15.18 ) 52.81 ( +1.05 ) 57.02 ( +1.96 ) 72.12 ( +15.48 ) 76.92 ( +15.84 )
Table 2: TTA comparisons.
Method CIFAR CUB Tiny
Zero-shot 61.07 51.76 56.64
Ensemble 63.66 51.54 61.39
TPT 64.09 51.83 62.77
C-TPT 64.86 52.54 63.20
Diff-TPT 63.04 52.80 61.60
VS2 64.52 52.69 58.14
VS2 + Ens 65.48 52.47 62.79
VS2 + Ens + Aug 65.80 51.54 62.59
Table 3: Inference-time compute.
Method GFLOPs GMACs Params (M)
Plain CLIP (vision) 8.7295 4.3623 87.456
LoRA (r=16r{=}16) 8.7885 4.3918 88.046
VS2 (SAE steering) 8.7342 4.3647 92.178
We evaluate VS2 on three datasets: CIFAR-100 (Krizhevsky et al., 2009), Tiny-ImageNet (Le and Yang, 2015), and CUB-200 (Wah et al., 2011), covering standard, complex, and fine-grained classification tasks. For the vision foundation model, we use CLIP (Radford et al., 2021b) with ViT-B/32, and ViT-B/16 backbones. Our goal is to investigate the effectiveness of Sparse Autoencoders (SAEs) for constructing steering vectors. In Sections 4.1 and 4.2, we focus on the case where the SAE can faithfully reconstruct the test inputs. To this end, we train top-$k$ SAEs on CLS token embeddings extracted from each layer of CLIP’s vision transformer, using the training split of each dataset to learn sparse latent representations (unsupervised in-domain data). Additional training details are provided in Appendix C. In Section 4.4, we relax this assumption by training a single, more general SAE across all train datasets, including both general and fine-grained tasks, to better approximate a real-world deployment scenario and evaluate how VS2 can avoid harmful steering when test samples cannot be faithfully reconstructed. This capability is novel compared to common steering vector and test-time adaptation techniques.
(a) CIFAR-100
(b) CUB-200
(c) Tiny-ImageNet
4.1 VS2: Visual Sparse Steering
@@
We evaluate VS2 on downstream image classification and compare against baselines that isolate reconstruction and latent amplification effects. We report: (1) the zero-shot CLIP model ($\text{CLIP}_{\text{ZS}}$); (2) $\text{SAE}_{\text{REC}}^{\text{F}}$, which replaces the final-layer [CLS] token with its SAE reconstruction (no steering); and (3) $\text{SAE}_{\text{REC}}^{\text{A}}$, which reconstructs [CLS] tokens from all layers (no steering). These baselines measure the impact of reconstruction alone. We additionally include $\text{SAE}_{\text{REC}}^{\text{F}+\gamma}$, which scales the top-$k$ latent activations by a fixed factor ($\gamma=1.5$) before reconstructing the final-layer [CLS] token, serving as an ablation for latent amplification without our additive steering direction. Appendix G provides pseudocode for these variants. We also report comparisons to Splice (Bhalla et al., 2024) in Appendix H, which defines latents using an external vocabulary rather than learning an unsupervised SAE dictionary. Unless stated otherwise, VS2 applies steering to the final-layer [CLS] token; Appendix E analyzes steering across layers.
@@ Results.
From Table 1, we show that using only SAE reconstructions typically reduces zero-shot performance; we attribute this degradation to reconstruction error in SAEs (Engels et al., 2025). However, when SAE features are amplified during reconstruction, performance can improve in some cases, as demonstrated by $\text{SAE}_{\text{REC}}^{\text{F}+\gamma}$ on CIFAR-100. Specifically, $\text{SAE}_{\text{REC}}^{\text{F}+\gamma}$ outperforms the zero-shot baseline on CIFAR-100 by 1.62% and 2.85% with ViT-B/32 and ViT-B/16, respectively, but underperforms CLIP on CUB-200 and Tiny-ImageNet. In contrast, VS2 consistently improves over zero-shot CLIP on all three datasets and both backbones, yielding gains of 3.45% and 4.12% on CIFAR-100, 0.93% and 1.08% on CUB-200, and 1.50% and 1.84% on Tiny-ImageNet for ViT-B/32 and ViT-B/16, respectively. These results demonstrate that sparsity-guided steering outperforms both reconstruction-only and amplification-only baselines. Although our focus is on how to steer ViTs using SAE features rather than interpretability, SAEs are often studied in mechanistic interpretability to identify human-interpretable features. Accordingly, Appendix A provides qualitative examples to assess whether the learned sparse features are consistent with subtle class-specific visual attributes (e.g., gray upperparts, white underparts, white throat, and black eyes in bird images). Finally, Figure 2 and Appendix B report a sensitivity analysis over the steering hyperparameters $\gamma$ and $\lambda$, showing a broad region of near-optimal performance and supporting the robustness of VS2 to reasonable hyperparameter choices.
Comparison with Test-Time Adaptation Baselines.
Table 2 compares VS2 to test-time prompt tuning methods TPT (Shu et al., 2022), C-TPT (Yoon et al., 2024), and Diff-TPT (Feng et al., 2023) using ViT-B/32. These baselines adapt by optimizing text prompts at test time, whereas VS2 performs no test-time optimization and does not update prompts or model weights. The Ensemble setting uses 80 fixed ImageNet-style templates following CLIP’s standard zero-shot protocol (Radford et al., 2021a). Aug denotes test-time image augmentations. The TTA baselines follow the standard 63-view protocol, while our VS2+Ens+Aug ablation uses 10 augmented views per image. Since the compared methods use different adaptation mechanisms and test-time compute, we treat this as an accuracy reference under common TTA evaluation protocols rather than a matched-compute comparison. Under this evaluation, VS2 matches or exceeds prompt-tuning baselines on CIFAR-100 and remains competitive on CUB-200; moreover, combining VS2 with prompt ensembling closes much of the gap on Tiny-ImageNet (VS2+Ens: 62.79% vs. TPT: 62.77%).
Computational Overhead.
VS2 is designed to be a lightweight alternative to parameter- or prompt-based test-time adaptation. Table 3 reports inference-time cost for the CLIP ViT-B/32 vision encoder under different adaptation strategies. Adding VS2 increases vision-encoder compute by only 0.0047 GFLOPs ($0.1\%$), since the SAE operates only on the [CLS] token at a single layer. In contrast, a LoRA adapter with rank 16 is applied inside every transformer block, leading to higher FLOPs despite fewer trainable parameters. At training time, learning the SAE is also inexpensive. With cached activations, SAE training requires only 0.014 GFLOPs per sample, compared to roughly 26 GFLOPs per sample for LoRA or full fine-tuning of the vision encoder; full training-time FLOPs and parameter counts are reported in Appendix M.
Finally, prompt-optimization-based TTA can be substantially more expensive at inference because it evaluates many augmented views and performs gradient-based updates. A single CLIP forward pass on a $224\times 224$ image with 100 text prompts costs approximately 0.59 TFLOPs (vision+text; Appendix M). Under the standard TPT protocol (63 augmented views and 4 optimization steps with a backward pass through the text encoder), this increases to roughly 9.8 TFLOPs per test image, i.e., $\approx 16.6\times$ the cost of plain CLIP (Appendix M). In contrast, VS2 adds only one SAE forward pass and keeps the per-image vision-encoder overhead below $0.1\%$ of CLIP.
4.2 VS2++: Upper-Bound Analysis via Selective Amplification
In VS2, we uniformly amplify the active sparse features of the test image. A natural extension is selective amplification: upweight features that are predictive for the target class while suppressing nuisance factors. To quantify the headroom available from accurate feature selection, we introduce VS2++ as a retrieval-based upper-bound analysis that uses neighbors to induce weak supervision for selecting which sparse features to steer.
Concretely, for each test image we retrieve the top-$50$ nearest neighbors from the dataset’s unlabeled training split using DINOv2 (Oquab et al., 2023). We then form pseudo-positive and pseudo-negative groups using CLIP predictions: neighbors with the most frequent pseudo-label constitute positives, and the remaining neighbors are treated as negatives. We construct a contrastive steering direction in SAE feature space by subtracting the average negative sparse activations from the average positive sparse activations. This corresponds to the non-oracle setting in Table 1. For comparison, we also report an oracle upper bound where positives and negatives are defined using the ground-truth labels of the retrieved neighbors.
We emphasize that VS2++ is used primarily to study the gains possible from reliable feature selection, motivating future work on robust selection under noisy pseudo-labels. As baselines, we include a weighted Retrieval Augmented Generation (RAG) pipeline (Lewis et al., 2020), which combines the query embedding with a weighted average of retrieved CLIP embeddings (Appendix D), as well as a non-SAE contrastive steering baseline that computes the difference between the mean CLIP embeddings of the positive and negative groups.